OpenStudy (anonymous):

Algebra 9th grade help please! A medal will be rewarded! :)

OpenStudy (anonymous):

@phi

OpenStudy (anonymous):

@Phoenix515

OpenStudy (anonymous):

Here's the question: Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 8^{x} and y = 2^{x + 2} intersect are the solutions of the equation 8^{x} = 2^{x + 2}. Part B Make tables to find the solution to 8^{x} = 2^{x + 2}. Take the integer values of x between -3 and 3. Part C: How can you solve the equation 8^{x} = 2^{x + 2} graphically.

OpenStudy (anonymous):

@e.mccormick may I please have some help? I am having a hard time w/ this!

OpenStudy (anonymous):

Anyone? antschauble I need some help.

OpenStudy (anonymous):

are you solving it

OpenStudy (e.mccormick):

Well, I am going to be leaving as soon as someone is ready.... but here is a graph to start with: https://www.desmos.com/calculator/i6iwc6vaud

OpenStudy (anonymous):

Could you please explain? I don't understand.

OpenStudy (e.mccormick):

Like I said, I don't have time. I am literally waiting for someone so we can walk out the door. Mathematically, here is a clue: \((a^m)^n=a^{m\cdot n}\) \(8=2^3\)

OpenStudy (e.mccormick):

With the math info there, you can use \(\log_2\) to solve it.

OpenStudy (anonymous):

ok, thank you!

OpenStudy (anonymous):

@phi may i have some help?

OpenStudy (anonymous):

@amistre64 may I have some help

OpenStudy (amistre64):

i have time constrictions as well. srry

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

@JM98SMITH may I have some help

OpenStudy (anonymous):

@jhonyy9

OpenStudy (phi):

which part did you do ?

OpenStudy (anonymous):

None of them, I don't know what to do, could you please explain? I mean, I'm confused as to what to do. :/

OpenStudy (phi):

Can you do Part B: Part B Make tables to find the solution to 8^{x} = 2^{x + 2}. Take the integer values of x between -3 and 3. for each x from -3 to +3, figure out 8^x and 2^(x+2) and put that into a table.

OpenStudy (phi):

make 3 columns, with the headings "x", "8^x" and "2^(x+2)"

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I understand, so make a column for x, w/ the numbers going from -3,-2,-1,0 up to +3, plug those numbers into the equations above, and then what do I do for part c.?

OpenStudy (phi):

Plot the points from part B, and connect the dots. you should get two curves that cross. where they cross is the answer.

OpenStudy (anonymous):

Thank you so much for the help.

OpenStudy (phi):

For Part A Part A: Explain why the x-coordinates of the points where the graphs of the equations y = 8^{x} and y = 2^{x + 2} intersect are the solutions of the equation 8^{x} = 2^{x + 2}. Let's say that the point (x0,y0) "works" for both equations. In other words, we get \[ 8^{x_0}= y_0 \text{ and } 2^{x_0+2} = y_0 \] The point is both equations equal the same number y0. That means \[ 8^{x_0}=2^{x_0+2} \] and the number x0 is the solution. Because the point (x0,y0) is "on" both curves, it is the intersection of the two curves. The x value of the intersection point will be the answer